6(t+3) + 8 - 5(t + 2)= can you tell me the steps to figure this out: 6(t+3) + 8 - 5(t + 2)= can you tell me the steps to figure this out
Answer by Fombitz(1755)

6(t+3)+8-5(t+2)=(6t+18)+8-(5t+10) First step, distribute 6 and 5 into parentheses.
6(t+3)+8-5(t+2)=6t+18+8-5t-10) Simplify.
6(t+3)+8-5(t+2)=(6t-5t)+(18+8-10) Group similar terms (t terms and constants).
6(t+3)+8-5(t+2)=t+16 Simplify.